# ShortV: Efficient Multimodal Large Language Models by Freezing Visual Tokens in Ineffective Layers

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) suffer from high computational costs due to their massive size and the large number of visual tokens. In this paper, we investigate layer-wise redundancy in MLLMs by introducing a novel metric, Layer Contribution (LC), which quantifies the impact of a layer’s transformations on visual and text tokens, respectively. The calculation of LC involves measuring the divergence in model output that results from removing the layer’s transformations on the specified tokens. Our pilot experiment reveals that many layers of MLLMs exhibit minimal contribution during the processing of visual tokens. Motivated by this observation, we propose ShortV, a training-free method that leverages LC to identify ineffective layers, and freezes visual token updates in these layers. Experiments show that ShortV can freeze visual token in approximately 60% of the MLLM layers, thereby dramatically reducing computational costs related to updating visual tokens. For example, it achieves a 50% reduction in FLOPs on LLaVA-NeXT-13B while maintaining superior performance. The code will be publicly available at https://github.com/icip-cas/ShortV.
@@ 3 ShortV
| Method | TFLOPs | FLOPs | VQAv2 | GQA | SEED- | MMMU | MME | MMBench | MMStar |
| Ratio | Bench | (val) | EN | ||||||
| LLaVA-1.5-7B | |||||||||
| Vanilla | 8.5 | 100% | 76.5 | 61.9 | 66.1 | 36.3 | 1510.7 | 64.1 | 33.7 |
| FastV (=2, =50%) | 4.9 | 58% | 73.5 | 60.2 | 65.4 | 35.8 | 1475.6 | 64.3 | 32.4 |
| VTW (=16) | 4.7 | 55% | 66.3 | 55.1 | 66.2 | 36.1 | 1497.0 | 64.0 | 32.8 |
| ShortV (Ours, =19) | 4.7 | 55% | 75.7 | 60.9 | 66.2 | 36.2 | 1503.1 | 64.8 | 33.3 |
| LLaVA-1.5-13B | |||||||||
| Vanilla | 16.6 | 100% | 78.0 | 63.3 | 68.2 | 35.4 | 1531.3 | 68.9 | 36.1 |
| FastV (=2, =50%) | 9.4 | 57% | 76.7 | 59.4 | 67.8 | 34.6 | 1506.6 | 68.3 | 35.9 |
| VTW (=20) | 9.1 | 55% | 75.3 | 60.6 | 68.2 | 34.9 | 1533.0 | 68.5 | 36.1 |
| ShortV (Ours, =24) | 9.1 | 55% | 77.2 | 62.0 | 68.0 | 35.8 | 1535.9 | 68.6 | 37.1 |
| LLaVA-NeXT-7B | |||||||||
| Vanilla | 42.7 | 100% | 80.0 | 64.1 | 70.2 | 36.4 | 1519.0 | 67.1 | 37.1 |
| FastV (=2, =50%) | 22.0 | 52% | 79.5 | 63.0 | 69.6 | 35.1 | 1482.0 | 66.3 | 36.5 |
| VTW (=16) | 21.8 | 51% | 75.6 | 55.8 | 70.2 | 35.7 | 1518.2 | 67.1 | 37.6 |
| ShortV (Ours, =19) | 21.6 | 51% | 78.8 | 63.4 | 70.4 | 36.0 | 1525.1 | 67.2 | 37.8 |
| LLaVA-NeXT-13B | |||||||||
| Vanilla | 81.8 | 100% | 80.9 | 65.7 | 71.9 | 35.9 | 1570.0 | 69.3 | 39.9 |
| FastV (=2, =50%) | 42.1 | 51% | 76.8 | 62.9 | 71.5 | 35.9 | 1546.4 | 68.5 | 39.6 |
| VTW (=20) | 41.7 | 51% | 77.0 | 61.5 | 71.8 | 34.8 | 1569.4 | 69.1 | 39.8 |
| ShortV (Ours, =24) | 41.0 | 50% | 79.7 | 63.6 | 71.8 | 36.2 | 1553.0 | 70.2 | 39.9 |
@@ 4.2 Main Results
In this section, we conduct experiments to compare ShortV with the baselines. The results are shown in Table 1. We provide the details for selecting replaced layers and their layer ids in Appendix B. We perform evaluation on multiple popular vision language benchmarks, including MME [15], MMBench [33], MMMU (val) [55], MMStar [9], SEED-Bench [25], VQAv2 [16], and GQA [19]. We manually choose the number of ShortV layers to maintain a similar or lower FLOPs ratio compared with the baselines. The FLOPs ratio of ShortV is calculated according to Equation 5, where we set the number of text tokens to 64. Specifically, we choose =19 for the 7B models, and =24 for the 13B models. As shown in Table 1, our ShortV achieves comparable or superior performance across multiple benchmarks compared with the baselines.
@@ 4.3 Balance between Efficiency and Performance
In this section, we conduct an experiment to investigate the impact of ShortV’s parameter , which denotes the number of ShortV layers. The experimental results are presented in Table 2. We also provide additional comprehensive results across more settings in Appendix D. To facilitate intuitive comprehension, we plot the correlation between the normalized average score on benchmarks and the number of ShortV layers in Figure 1(b).

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) suffer from high computational costs due to their massive size and the large number of visual tokens. In this paper, we investigate layer-wise redundancy in MLLMs by introducing a novel metric, Layer Contribution (LC), which quantifies the impact of a layer’s transformations on visual and text tokens, respectively. The calculation of LC involves measuring the divergence in model output that results from removing the layer’s transformations on the specified tokens. Our pilot experiment reveals that many layers of MLLMs exhibit minimal contribution during the processing of visual tokens. Motivated by this observation, we propose ShortV, a training-free method that leverages LC to identify ineffective layers, and freezes visual token updates in these layers. Experiments show that ShortV can freeze visual token in approximately 60% of the MLLM layers, thereby dramatically reducing computational costs related to updating visual tokens. For example, it achieves a 50% reduction in FLOPs on LLaVA-NeXT-13B while maintaining superior performance. The code is publicly available at https://github.com/icip-cas/ShortV.
@@ 3 ShortV
| Method | TFLOPs | FLOPs | VQAv2 | GQA | SEED- | MMMU | MME | MMBench | MMStar | Flickr30K |
| Ratio | Bench | (val) | EN | CIDEr | ||||||
| LLaVA-1.5-7B | ||||||||||
| Vanilla | 8.5 | 100% | 76.5 | 61.9 | 66.1 | 36.3 | 1510.7 | 64.1 | 33.7 | 74.9 |
| FastV (=2, =50%) | 4.9 | 58% | 73.5 | 60.2 | 65.4 | 35.8 | 1475.6 | 64.3 | 32.4 | 67.5 |
| VTW (=16) | 4.7 | 55% | 66.3 | 55.1 | 66.2 | 36.1 | 1497.0 | 64.0 | 32.8 | 44.5 |
| ShortV (Ours, =19) | 4.7 | 55% | 75.7 | 60.9 | 66.2 | 36.2 | 1503.1 | 64.8 | 33.3 | 71.3 |
| LLaVA-1.5-13B | ||||||||||
| Vanilla | 16.6 | 100% | 78.0 | 63.3 | 68.2 | 35.4 | 1531.3 | 68.9 | 36.1 | 79.6 |
| FastV (=2, =50%) | 9.4 | 57% | 76.7 | 59.4 | 67.8 | 34.6 | 1506.6 | 68.3 | 35.9 | 73.4 |
| VTW (=20) | 9.1 | 55% | 75.3 | 60.6 | 68.2 | 34.9 | 1533.0 | 68.5 | 36.1 | 65.9 |
| ShortV (Ours, =24) | 9.1 | 55% | 77.2 | 62.0 | 68.0 | 35.8 | 1535.9 | 68.6 | 37.1 | 76.4 |
| LLaVA-NeXT-7B | ||||||||||
| Vanilla | 42.7 | 100% | 80.0 | 64.1 | 70.2 | 36.4 | 1519.0 | 67.1 | 37.1 | 69.7 |
| FastV (=2, =50%) | 22.0 | 52% | 79.5 | 63.0 | 69.6 | 35.1 | 1482.0 | 66.3 | 36.5 | 66.7 |
| VTW (=16) | 21.8 | 51% | 75.6 | 55.8 | 70.2 | 35.7 | 1518.2 | 67.1 | 37.6 | 38.4 |
| ShortV (Ours, =19) | 21.6 | 51% | 78.8 | 63.4 | 70.4 | 36.0 | 1525.1 | 67.2 | 37.8 | 65.7 |
| LLaVA-NeXT-13B | ||||||||||
| Vanilla | 81.8 | 100% | 80.9 | 65.7 | 71.9 | 35.9 | 1570.0 | 69.3 | 39.9 | 66.7 |
| FastV (=2, =50%) | 42.1 | 51% | 76.8 | 62.9 | 71.5 | 35.9 | 1546.4 | 68.5 | 39.6 | 66.0 |
| VTW (=20) | 41.7 | 51% | 77.0 | 61.5 | 71.8 | 34.8 | 1569.4 | 69.1 | 39.8 | 56.6 |
| ShortV (Ours, =24) | 41.0 | 50% | 79.7 | 63.6 | 71.8 | 36.2 | 1553.0 | 70.2 | 39.9 | 67.5 |
@@ 4.2 Main Results
In this section, we conduct experiments to compare ShortV with the baselines. The results are shown in Table 1. We perform evaluation on multiple popular vision language benchmarks, including MME [15], MMBench [33], MMMU (val) [55], MMStar [9], SEED-Bench [25], VQAv2 [16], and GQA [19]. We manually choose the number of ShortV layers to maintain a similar or lower FLOPs ratio compared with the baselines. The FLOPs ratio of ShortV is calculated according to Equation 5, where we set the number of text tokens to 64. Specifically, we choose =19 for the 7B models, and =24 for the 13B models. As shown in Table 1, our ShortV achieves comparable or superior performance across multiple benchmarks compared with the baselines.
@@ 4.3 Balance between Efficiency and Performance
In this section, we conduct an experiment to investigate the impact of ShortV’s parameter , which denotes the number of ShortV layers. The experimental results are presented in Table 2. To facilitate intuitive comprehension, we plot the correlation between the normalized average score on benchmarks and the number of ShortV layers in Figure 1(b).
@@ 6 Conclusion
7 Acknowledgment
We sincerely thank the reviewers for their insightful comments and valuable suggestions. This work was supported by Beijing Natural Science Foundation (L243006), Beijing Municipal Science and Technology Project (Nos. Z231100010323002), the Natural Science Foundation of China (No. 62306303, 62476265) and the Basic Research Program of ISCAS (Grant No. ISCAS-ZD-202401).
@@ Appendix D ShortV under Different Settings
![[Uncaptioned image]](fig/cartoon_image.jpg)
![[Uncaptioned image]](fig/fig_tricky_waikiki.jpg)